OpenStudy (anonymous):

sin^2(t) + tan^2(t)+cos^2(t) equals:

OpenStudy (anonymous):

\(sec^2 \theta\)

OpenStudy (anonymous):

You just have to know some trig identities: sin^2(t) + cos^(t) = 1 and then 1 + tan^2(t) = sec^2(t) Enjoy!

OpenStudy (anonymous):

Use: \[\sin^2 \theta + \cos^2 \theta = 1\] \[1 + \tan^2 \theta = \sec^2 \theta\]

OpenStudy (anonymous):

Can you show me how you got that?

OpenStudy (zzr0ck3r):

memorise them

OpenStudy (anonymous):

See: Rearrange the terms: \[(\sin^2 \theta + \cos^2 \theta) + \tan^2 \theta\] Use first identity here..

OpenStudy (anonymous):

I got it, thank you!

OpenStudy (zzr0ck3r):

or derive them, cos^2(x)+sin^2(x) = 1 devide all by cos^(x) and see what happens or devide them all by sin^2(x) and see what happens...

OpenStudy (anonymous):

Welcome dear..